# Bridging doubly heavy tetraquark mass spectrum with heavy baryons utilizing heavy antiquark-diquark symmetry

### v1

@@ Abstract
Motivated by the observation of the doubly charmed tetraquark $T_{cc}(3875)^{+}$, we present a systematic study of double heavy tetraquarks ($T_{QQ^{\prime}\bar{q}\bar{q}^{\prime}}$) using heavy antiquark-diquark symmetry (HADS) within a constituent quark model. By calibrating model parameters to known hadron spectra and incorporating the effective mass formula, we predict the masses for 38 ground-state tetraquarks with $cc$, bbbb, and $bc$ heavy quark pairs, including the non-strange, single-strange, and double-strange configurations with quantum numbers $J^{P}=0^{+},1^{+}$ and $2^{+}$. Notably, we identify several stable states below the relevant meson-meson thresholds, particularly in the $bb\bar{q}\bar{q}^{\prime}$ sector. The explicit connection between doubly heavy tetraquark and heavy baryon spectra through HADS reduces model dependence and reveals fundamental systematics in the heavy-quark hadron landscape.
@@ I Introduction
The discovery of exotic hadrons beyond the conventional quark-antiquark (meson) and three-quark (baryon) paradigms stands as a pivotal achievement of modern particle physics. Since 2003, many exotic hadronic states which cannot be well understood by the conventional constituent quark model have been observed. Typically, in the heavy flavor sector, the observations of the so-called $X,Y,Z$ states, the tetraquarks, the pentaquarks and so on, have extended the hadron spectra and deepen our understanding on the strong interactions between quarks in the nonperturbative region (see, e.g., Refs. Hosaka et al. (2016); Ali et al. (2017); Guo et al. (2018); Esposito et al. (2017); Lebed et al. (2017); Richard (2016); Chen et al. (2016); Liu et al. (2019); Brambilla et al. (2020); Liu et al. (2025) for reviews).
@@
In 2017, the doubly heavy baryon $\Xi_{cc}^{++}$ was discovered in the $\Lambda_{c}^{+}K^{-}\pi^{+}\pi^{+}$ mass spectrum by the LHCb collaboration with a mass $3621.4$ MeV Aaij et al. (2017), which is the first observed doubly heavy hadron ever. In 2021, the observation of the doubly charmed tetraquark candidate $T_{cc}^{+}(3875)$ by the LHCb collaboration marks a highlight moment Aaij et al. (2022, 2021). Its striking proximity to the $D^{*+}D^{0}$ mass threshold (within $\sim$360 keV) and narrow width ignited intense theoretical debate concerning its fundamental nature. Two primary interpretations emerged: a loosely bound hadronic molecule held together by residual strong forces between $D^{*}$ and $D$ mesons Du et al. (2022); Padmanath and Prelovsek (2022); Feijoo et al. (2021); Meng et al. (2021); Deng and Zhu (2022), and a compact tetraquark state with internal structure dominated by strong, short-distance correlations within the $c\bar{c}u\bar{d}$ valence quarks Karliner and Rosner (2017); Agaev et al. (2022); Kim et al. (2022); Wu and Ma (2023). All in all, studies so far have not determined the internal structure and nature of $T_{cc}(3875)^{+}$, the doubly heavy tetraquarks needs more studies to reveal its nature. While both pictures can qualitatively accommodate the mass and quantum numbers of the $T_{cc}^{+}(3875)$, they predict significantly different properties for its partners and the broader spectrum of doubly heavy tetraquarks (DHTs). Unambiguously resolving this structural dichotomy remains a central challenge and arouse the interests of studies on the doubly heavy hadrons both theoretically and experimentally Karliner and Rosner (2017); Olamaei et al. (2020); Dong et al. (2021); Chen et al. (2021); Ren et al. (2022); Deng and Zhu (2022); Xin and Wang (2022); Albaladejo (2022); Ling et al. (2022); Meng et al. (2021); Fleming et al. (2021); Huang et al. (2021); Du et al. (2022); Feijoo et al. (2021); Wu et al. (2022); Luo et al. (2022); Yan and Valderrama (2022); Padmanath and Prelovsek (2022); Agaev et al. (2022); Kim et al. (2022); Wu and Ma (2023); Ebert et al. (2007); Luo et al. (2017); Lü et al. (2020); Zhang et al. (2021); Weng et al. (2022); Mutuk (2024); Ma et al. (2024).
@@
Current theoretical efforts to map the DHTs landscape employ diverse methodologies. In Ref. Karliner and Rosner (2017), the authors have successfully predicted the existence of the doubly heavy tetraquarks before the discovery of $T_{cc}(3875)^{+}$ employing the constituent quark model with a good light quark symmetry. The masses of tetraquarks were calculated in the framework of the diquark-antidiquark picture in the relativistic quark model Ebert et al. (2007), in a relativized quark model within the variational method Lü et al. (2020), and an extended chromomagnetic model Weng et al. (2022). Ref. Luo et al. (2017) systematically studied the mass splittings of the DHT states with the color-magnetic interaction by considering color mixing effects and estimate roughly their masses. The masses and magnetic moments of tetraquarks were computed in a unified framework of the MIT bag model Zhang et al. (2021) and diffusion Monte Carlo method (DMC) Mutuk (2024). In Ref. Ma et al. (2024), the authors used DMC to calculate the DHT system in two kinds of constituent quark models, the pure constituent quark model AL1/AP1 and the chiral constituent quark model.
@@ II Theoretical framework
HADS is a significant concept in the study of hadronic physics Savage and Wise (1990); Georgi and Wise (1990); Carone (1991). In the limit of heavy quark masses approaching infinity in Quantum Chromodynamics (QCD), this symmetry emerges, connecting the properties of hadrons with two heavy quarks $QQ^{\prime}$ to those with one heavy antiquark $\bar{Q}$. Physically, HADS arises when a diquark pair forms a tightly bound, nearly point-like object. For instance, the attraction between two heavy quarks in a diquark comes from a color Coulombic interaction in the color $3_{c}$ channel. When the quark masses are large enough, the heavy quarks move slowly, acting like non-relativistic particle sources in a Coulombic potential. As the size of a Coulombic bound state is inversely proportional to its mass (for a fixed coupling), in the large mass limit, the diquark becomes a heavy, small object with color $\bar{3}_{c}$, effectively acting as a static point-like $\bar{3}_{c}$ color source, similar to a heavy antiquark.
This symmetry has important implications for understanding the spectra and properties of hadrons. It relates the quark structure of doubly heavy baryons to that of heavy antimesons. For example, the masses of $\Xi_{cc}^{(*)}$ baryons and $D^{(*)}$ mesons serves the relation $m_{\Xi_{cc}^{*}}-m_{\Xi_{cc}}=\frac{3}{4}(m_{\bar{D}^{*}}-m_{\bar{D}})$, which has been numerically confirmed by the experiments and lattice QCD simulations Padmanath et al. (2015); Chen and Chiu (2017); Alexandrou and Kallidonis (2017); Mathur and Padmanath (2019). Along this line, the masses and configurations of DHTs can also be related to those of heavy baryons by leveraging HADS, which relates doubly heavy tetraquarks to heavy baryons by treating the double-heavy diquark ($QQ^{\prime}$) as a heavy antiquark $\bar{X}$. Combined with the constituent quark model—well-established for describing meson and baryon spectra—this symmetry allows us to predict the mass spectrum of doubly heavy tetraquarks systematically. We extend previous studies to include not only doubly charmed ($T_{cc}$) and doubly bottomed ($T_{bb}$) states but also bottom-charmed ($T_{bc}$) configurations, covering non-strange, single-strange, and double-strange light quark combinations. The predicted states span spin-parity quantum numbers $0^{+}$, $1^{+}$ and $2^{+}$, providing a complete landscape of ground states of DHTs.
@@ II.1 Configurations of DHTs
The configurations of DHTs are labeled with $\bar{q}\bar{q}^{\prime}]^{c}_{s}[QQ^{\prime}]^{C}_{S}]^{I}_{J}$. Where $C$ and $S$ is the color and spin of the heavy diquark $QQ^{\prime}$, respectively. $c$ and $s$ is the color and spin of the light diquark $\bar{q}\bar{q}^{\prime}$, respectively. $I$ is the isospin of the tetraquark and $J$ is the total angular momentum of the tetraquark. Doubly heavy tetraquarks with different number of strange quarks have different configurations. Since we are establishing a connection between DHTs and heavy baryons, we assign $c$ and $C$ to be $3$ and $\bar{3}$, respectively. For $s$ and $S$, their values can be either 0 or 1, depending on the flavor combination of the tetraquark. The configurations are listed below with three combinations on order of numbers of strange quarks, unflavored light quarks are denoted as $n$ with $n=u,d$.
@@ II.2 Mass formula of DHT in CQM
where $m_{i},B_{ij}$ and $\alpha_{ij}$ are $\cdots$, respectively. $F_{i},\sigma_{i}$ are the color and spin operators with $i,j=1,2,3$, respectively. This simple formula yields an average mass difference $\chi_{M}=7.6$ MeV. The method is successfully extended to doubly charmed and doubly bottomed tetraquarks with non-strange light quarks Wu and Ma (2023). Here in this work, besides what is studied before, we extend the study to bottom-charmed sector and cases that with non-, single-, and double-strange quark(s). Thus, the complete mass spectrum of the whole ground states of DHTs correlated to the heavy baryons are displayed.
@@
$\begin{split}m_{[QQ^{\prime}]^{C}_{S}}=&m_{Q}+m_{Q^{\prime}}+F_{Q}\cdot F_{Q^{\prime}}[B_{QQ^{\prime}}+\frac{\alpha_{Q_{1}Q_{2}}\sigma_{Q}\cdot\sigma_{Q^{\prime}}}{m_{Q}m_{Q^{\prime}}}].\end{split}$ (6)
@@
where $m_{Q^{(\prime)}}$ is the heavy quark mass. The quark masses, bindings and spin hyperfine coupling constants are needed to calculate the mass of the heavy diquark $QQ^{\prime}]^{C}_{S}$. All the required parameters are listed in Table 1, of which the values are determined by fitting the baryon spectra and corresponding hidden-flavor and bottom-charmed heavy mesons Wu and Ma (2023). The masses of the heavy diquarks are determined to be $m_{[cc]^{\bar{3}}_{1}}=2m^{b}_{c}+(-2/3)(B_{cc}+\alpha_{cc}/(m^{b}_{c})^{2})=3300.8$ MeV, $m_{[bb]^{\bar{3}}_{1}}=2m^{b}_{b}+(-2/3)(B_{bb}+\alpha_{bb}/(m^{b}_{b})^{2})=9821.0$ MeV, $m_{[cb]^{\bar{3}}_{0}}=m^{b}_{c}+m^{b}_{b}+(-2/3)(B_{cb}-3\alpha_{cb}/(m^{b}_{b}\cdot m^{b}_{c}))=6567.0$ MeV and $m_{[cb]^{\bar{3}}_{1}}=m^{b}_{c}+m^{b}_{b}+(-2/3)(B_{cb}+\alpha_{cb}/(m^{b}_{b}\cdot m^{b}_{c}))=6600.9$ MeV, respectively.

### v2

@@ Abstract
Motivated by the observation of the doubly charmed tetraquark $T_{cc}(3875)^{+}$, we present a systematic study of doubly heavy tetraquarks ($T_{QQ^{\prime}\bar{q}\bar{q}^{\prime}}$) using heavy antiquark-diquark symmetry (HADS) within a constituent quark model. By calibrating model parameters to known hadron spectra and incorporating the effective mass formula, we predict the masses for 38 ground-state tetraquarks with $cc$, bbbb, and $bc$ heavy quark pairs, including the non-strange, single-strange, and double-strange configurations with quantum numbers $J^{P}=0^{+},1^{+}$ and $2^{+}$. Notably, we identify several stable states below the relevant meson-meson thresholds, particularly in the $bb\bar{q}\bar{q}^{\prime}$ sector. The explicit connection between doubly heavy tetraquark and heavy baryon spectra through HADS reduces model dependence and reveals fundamental systematics in the heavy-quark hadron landscape.
@@ I Introduction
The discovery of exotic hadrons beyond the conventional quark-antiquark (meson) and three-quark (baryon) paradigms stands as a pivotal achievement of modern particle physics. Since 2003, many exotic hadronic states which cannot be well understood by the conventional constituent quark model have been observed. Typically, in the heavy flavor sector, the observations of the so-called $X,Y,Z$ states, the tetraquarks, the pentaquarks and so on, have extended the hadron spectra and deepen our understanding on the strong interactions between quarks in the nonperturbative region (see, e.g., Refs. Hosaka et al. (2016); Ali et al. (2017); Guo et al. (2018); Esposito et al. (2017); Lebed et al. (2017); Richard (2016); Chen et al. (2016); Liu et al. (2019); Brambilla et al. (2020); Liu et al. (2025b); Wang (2025) for reviews).
@@
In 2017, the doubly heavy baryon $\Xi_{cc}^{++}$ was discovered in the $\Lambda_{c}^{+}K^{-}\pi^{+}\pi^{+}$ mass spectrum by the LHCb collaboration with a mass $3621.4$ MeV Aaij and others (2017), which is the first observed doubly heavy hadron ever. In 2021, the observation of the doubly charmed tetraquark candidate $T_{cc}^{+}(3875)$ by the LHCb collaboration marks a highlight moment Aaij and others (2022, 2021). Its striking proximity to the $D^{*+}D^{0}$ mass threshold (within $\sim$360 keV) and narrow width ignited intense theoretical debate concerning its fundamental nature. Two primary interpretations emerged: a loosely bound hadronic molecule held together by residual strong forces between $D^{*}$ and $D$ mesons Du et al. (2022); Padmanath and Prelovsek (2022); Feijoo et al. (2021); Meng et al. (2021); Deng and Zhu (2022), and a compact tetraquark state with internal structure dominated by strong, short-distance correlations within the $c\bar{c}u\bar{d}$ valence quarks Karliner and Rosner (2017); Agaev et al. (2022); Kim et al. (2022); Wu and Ma (2023). All in all, studies so far have not determined the internal structure and nature of $T_{cc}(3875)^{+}$, the doubly heavy tetraquarks needs more studies to reveal its nature. While both pictures can qualitatively accommodate the mass and quantum numbers of the $T_{cc}^{+}(3875)$, they predict significantly different properties for its partners and the broader spectrum of doubly heavy tetraquarks (DHTs). Unambiguously resolving this structural dichotomy remains a central challenge and arouse the interests of studies on the doubly heavy hadrons both theoretically and experimentally Karliner and Rosner (2017); Eichten and Quigg (2017); Wang and Yan (2018); Wang (2018); Cheng et al. (2021); Chen et al. (2021); Ren et al. (2022); Deng and Zhu (2022); Xin and Wang (2022); Albaladejo (2022); Ling et al. (2022); Meng et al. (2021); Fleming et al. (2021); Feijoo et al. (2021); Guo et al. (2022); Du et al. (2022); Huang et al. (2021); Wu et al. (2022); Luo et al. (2022); Abreu (2022); Yan and Valderrama (2022); Padmanath and Prelovsek (2022); Agaev et al. (2022); Kim et al. (2022); Wu and Ma (2023); Ebert et al. (2007); Luo et al. (2017); Lü et al. (2020); Zhang et al. (2021); Weng et al. (2022); Mutuk (2024); Ma et al. (2024); Liu et al. (2025a).
@@
Current theoretical efforts to map the DHTs landscape employ diverse methodologies. In Ref. Karliner and Rosner (2017), the authors have successfully predicted the existence of the doubly heavy tetraquarks before the discovery of $T_{cc}(3875)^{+}$ employing the constituent quark model with a good light quark symmetry. The masses of tetraquarks were calculated in the framework of the diquark-antidiquark picture in the relativistic quark model Ebert et al. (2007), in a relativized quark model within the variational method  et al. (2020), and an extended chromomagnetic model Weng et al. (2022). Ref. Luo et al. (2017); Guo et al. (2022) systematically studied the mass splittings of the DHT states with the color-magnetic interaction by considering color mixing effects and estimate roughly their masses. The masses and magnetic moments of tetraquarks were computed in a unified framework of the MIT bag model Zhang et al. (2021) and diffusion Monte Carlo method (DMC) Mutuk (2024). In Ref. Ma et al. (2024), the authors used DMC to calculate the DHT system in two kinds of constituent quark models, the pure constituent quark model AL1/AP1 and the chiral constituent quark model.
@@ II Theoretical framework
HADS is a significant concept in the study of hadronic physics Savage and Wise (1990); Georgi and Wise (1990); Carone (1991); Wu et al. (2020). In the limit of heavy quark masses approaching infinity in Quantum Chromodynamics (QCD), this symmetry emerges, connecting the properties of hadrons with two heavy quarks $QQ^{\prime}$ to those with one heavy antiquark $\bar{Q}$. Physically, HADS arises when a diquark pair forms a tightly bound, nearly point-like object. For instance, the attraction between two heavy quarks in a diquark comes from a color Coulombic interaction in the color $3_{c}$ channel. When the quark masses are large enough, the heavy quarks move slowly, acting like non-relativistic particle sources in a Coulombic potential. As the size of a Coulombic bound state is inversely proportional to its mass (for a fixed coupling), in the large mass limit, the diquark becomes a heavy, small object with color $\bar{3}_{c}$, effectively acting as a static point-like $\bar{3}_{c}$ color source, similar to a heavy antiquark.
This symmetry has important implications for understanding the spectra and properties of hadrons. It relates the quark structure of doubly heavy baryons to that of heavy antimesons. For example, the masses of $\Xi_{cc}^{(*)}$ baryons and $D^{(*)}$ mesons serves the relation $m_{\Xi^{*}_{cc}}-m_{\Xi_{cc}}=\frac{3}{4}(m_{\bar{D}^{*}}-m_{\bar{D}})$ has been numerically confirmed with the masses observed by experiments ($m_{\Xi_{cc}}$, $m_{\bar{D}^{*}}$ and $m_{\bar{D}}$) and lattice QCD simulations ($m_{\Xi_{cc}^{*}}$Padmanath et al. (2015); Chen and Chiu (2017); Alexandrou and Kallidonis (2017); Mathur and Padmanath (2019). Along this line, the masses and configurations of DHTs can also be related to those of heavy baryons by leveraging HADS, which relates DHTs to heavy baryons by treating the double-heavy diquark ($QQ^{\prime}$) as a heavy antiquark $\bar{X}$. Combined with the constituent quark model—well-established for describing meson and baryon spectra—this symmetry allows us to predict the mass spectrum of DHTs systematically. We extend previous studies to include not only doubly charmed ($T_{cc}$) and doubly bottomed ($T_{bb}$) states but also bottom-charmed ($T_{bc}$) configurations, covering non-strange, single-strange, and double-strange light quark combinations. The predicted states span spin-parity quantum numbers $0^{+}$, $1^{+}$ and $2^{+}$, providing a complete landscape of ground states of DHTs.
@@ II.1 Configurations of DHTs
The configurations of DHTs are labeled with $\bar{q}\bar{q}^{\prime}]^{c}_{s}[QQ^{\prime}]^{C}_{S}]^{I}_{J}$. Where $C$ and $S$ is the color and spin of the heavy diquark $QQ^{\prime}$, respectively. $c$ and $s$ is the color and spin of the light diquark $\bar{q}\bar{q}^{\prime}$, respectively. $I$ is the isospin of the tetraquark and $J$ is the total angular momentum of the tetraquark. DHTs with different number of strange quarks have different configurations. Since we are establishing a connection between DHTs and heavy baryons, we assign $c$ and $C$ to be $3$ and $\bar{3}$, respectively. Color sextet combinations may also exist in the color structure of DHT states. However, such configurations have not been experimentally observed to date. In diquarks, the Casimir coefficient for the $\mathbf{3}_{c}\otimes\mathbf{3}_{c}=\mathbf{6}_{c}$ representation is $1/3$, indicating a repulsive interaction between the two quarks in this color configuration, while the coefficient for $\mathbf{3}_{c}\otimes\mathbf{3}_{c}=\bar{\mathbf{3}}_{c}$ is 2/3-2/3, suggesting a strong attractive interaction. This implies that $\mathbf{3}_{c}\otimes\mathbf{3}_{c}=\bar{\mathbf{3}}_{c}$ is likely the dominant form of diquark coupling. Studies in Ref.  et al. (2020) show that the weight of the latter configuration in DHT states exceeds 90%. Based on these theoretical and experimental reasons, in this work, we consider only the $\mathbf{3}_{c}$ and $\bar{\mathbf{3}}_{c}$ color configurations for diquarks. For $s$ and $S$, their values can be either 0 or 1, depending on the flavor combination of the tetraquark. The configurations are listed below with three combinations on order of numbers of strange quarks, unflavored light quarks are denoted as $n$ with $n=u,d$.
@@ II.2 Mass formula of DHT in CQM
where $m_{i},B_{ij}$ and $\alpha_{ij}$ are the mass of the $i$th constituent quark, the binding and hyperfine spin coupling constant between the $i$th and $j$th quarks, respectively. $F_{i},\sigma_{i}$ are the color and spin operators with $i,j=1,2,3$, respectively. This simple formula yields a standard mass variance $\chi_{M}=7.6$ MeV. The method is successfully extended to doubly charmed and doubly bottomed tetraquarks with non-strange light quarks Wu and Ma (2023). Here in this work, besides what is studied before, we extend the study to bottom-charmed sector and cases that with non-, single-, and double-strange quark(s). Thus, the complete mass spectrum of the whole ground states of DHTs correlated to the heavy baryons are displayed.
@@
$m_{[QQ^{\prime}]^{C}_{S}}=m_{Q}+m_{Q^{\prime}}+F_{Q}\cdot F_{Q^{\prime}}[B_{QQ^{\prime}}+\frac{\alpha_{QQ^{\prime}}\sigma_{Q}\cdot\sigma_{Q^{\prime}}}{m_{Q}m_{Q^{\prime}}}],$ (3)
@@
where $m_{Q^{(\prime)}}$ is the heavy quark mass. The quark masses, bindings and spin hyperfine coupling constants are needed to calculate the mass of the heavy diquark $QQ^{\prime}]^{C}_{S}$.
All the required parameters are listed in Table 1, of which the values are determined by fitting the baryon spectra and corresponding hidden-flavor and bottom-charmed heavy mesons Wu and Ma (2023); Karliner and Rosner (2014). It should be noted that the masses of constituent quarks in baryons are diffrent with those in mesons. In this work, since quarks are in the form of diquarks, the quark masses we adopt for calculating the mass spectra are all based on the quark masses within baryons. Only when determining the hyperfine splitting constants $\alpha_{QQ^{\prime}}$ and binding energies $B_{QQ^{\prime}}$ between heavy quarks, given that the corresponding heavy baryons have not yet been experimentally observed, we resort to meson masses for such determinations. Take $bb$ as example, since $m_{\Upsilon_{b}}=2m_{b}^{m}+(-4/3)[B^{m}_{bb}+\alpha_{bb}/(m_{b}^{m})^{2}]$ and $m_{\eta_{b}}=2m_{b}^{m}+(-4/3)[B^{m}_{bb}-3\alpha_{bb}/(m_{b}^{m})^{2}]$, we have $\alpha_{bb}/(m_{b}^{b})^{2}\approx\alpha_{bb}/(m_{b}^{m})^{2}=-3/16(m_{\Upsilon}-m_{\eta_{b}})=-11.6$ MeV and $B_{bb}\approx B^{m}_{bb}=-3/16[(3m_{\Upsilon_{b}}+m_{\eta_{b}})-8m^{m}_{b}]=422.0$ MeV, neglecting the small differences of constituent quark masses between baryons and mesons. The meson masses are taken from experiments and the heavy quark mass in mesons $m_{b}^{m}=5003.8$ MeV and $m_{b}^{m}=1663.2$ MeV (the superscript mm/$b$ denotes in mesons/baryons) are from Ref. Karliner and Rosner (2014) within the same model. For $B_{cb}$ and $\alpha_{cb}$, since $B^{*}_{c}$ is not observed yet, we have to refer to the theoretical predictions. In this work, we use the value in Ref. Karliner and Rosner (2014), $m_{B^{*}_{c}}-m_{B_{c}}=68$ MeV, which is determined by an interpolation method with the mass differences $m_{\Upsilon_{b}}-m_{\eta_{b}}=62.3$ MeV and $m_{J/\psi}-m_{\eta_{c}}=113.2$ MeV. The four heavy meson masses are from experiments and the interpolation method used are based on the same constituent quark model as in this work. Other theoretical predictions of $m_{B^{*}_{c}}-m_{B_{c}}$ are also alternative, for example, the lattice prediction yields a value of $55(3)$ MeV Mathur et al. (2018). Using the former value gives $B_{bc}=256.2\ \text{MeV}$ and $\alpha_{cb}/(m_{b}m_{c})=12.7$ MeV, while the latter value yields $B_{bc}=263.5\ \text{MeV}$ and $\alpha_{cb}/(m_{b}m_{c})=10.3$ MeV. This results in a shift of less than 10MeV10\ $\text${MeV} in the finally obtained $T_{bc}$ mass spectrum, which does not affect the main results and conclusions of this paper. We adopt the former value $m_{B^{*}_{c}}-m_{B_{c}}=68$ MeV for two reasons: one is based on the consideration of model consistency, and the other is that according to the heavy quark spin symmetry, the spin mass splitting of $bc$ mesons $m_{B^{*}_{c}}-m_{B_{c}}$ should lie between that of $cc$ and $bb$ mesons.